OpenStudy (anonymous):

x2 + 18x + 76 = 0 Which of the following is the first correct step to write the above equation in the form (x - p)2 = q, where p and q are integers? Add 9 to both sides of the equation Add 5 to both sides of the equation Subtract 5 from both sides of the equation Subtract 9 from both sides of the equation

OpenStudy (anonymous):

@jim_thompson5910 @zepdrix

OpenStudy (anonymous):

@Awesome781 @bribri50 @Compassionate @d3th21 @FBland1

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

can you help? @zzr0ck3r

OpenStudy (anonymous):

@ganeshie8 @Hero @jim_thompson5910

OpenStudy (anonymous):

@thomaster @Luigi0210

OpenStudy (anonymous):

@dan815 @iambatman @phi @jdoe0001

OpenStudy (anonymous):

@nincompoop @myininaya

jimthompson5910 (jim_thompson5910):

What's the middle coefficient?

OpenStudy (anonymous):

i thought it was 18x ... @jim_thompson5910

OpenStudy (anonymous):

or is it 9? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

18x is the middle term so 18 is the middle coefficient

jimthompson5910 (jim_thompson5910):

cut that in half to get 18/2 = 9 then square it to get 9^2 = ???

OpenStudy (anonymous):

so I would need to subtract 9 from both sides?

OpenStudy (anonymous):

that would be 81

jimthompson5910 (jim_thompson5910):

x^2 + 18x + 81 is a perfect square we have x^2 + 18x + 76

jimthompson5910 (jim_thompson5910):

x^2 + 18x + 76 = 0 x^2 + 18x + 76 + ____ = 0 + ____ x^2 + 18x + 81 = ____

OpenStudy (anonymous):

so I would subtract 5 from both sides?

jimthompson5910 (jim_thompson5910):

76 -5 = 71 not 81

OpenStudy (anonymous):

then adding it is. :)

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thanks for your help. I really appreciate it. :)

jimthompson5910 (jim_thompson5910):

you're welcome